OpenStudy (anonymous):

Monochromatic light of 605 nm falls on a single slit of width 0.095 mm. The slit is located 85 cm from a screen. How far from the center is the first dark band?

OpenStudy (anonymous):

\[y = ( m + 1/2) \lambda L/d\]

OpenStudy (anonymous):

Here... m=1 for first dark fringe.. L is the distance of the screen from slit

OpenStudy (anonymous):

d is the width of the slit, and lambda is the wavelength.

OpenStudy (anonymous):

You can use the same equation as for two slids, which is \[\frac{\Delta x}{d}=\frac{k \cdot \lambda}{g}\] \[ \Delta x \] is what you're looking for, lamda is the wavelength, g the width of your slit and d the distance between slit and screen. The difference to two or more slits is, that u dont find the bright, but the dark bands.

OpenStudy (anonymous):

\[y = ( 1 + 0.5)(605 \times 10^{-9} )(0.85) / ( 0.095 \times 10^{-3}\]

OpenStudy (anonymous):

\[Y = 8.11 \times 10^{-3}\]

OpenStudy (anonymous):

I came up with 0.008119737

OpenStudy (anonymous):

If it helps its on of these 4. I just need to understand these questions better its on my review sheet. 0.5 mm 5.4 mm 540 mm 5,400 mm

OpenStudy (anonymous):

and, as a proper physicist, don't forget your units

OpenStudy (anonymous):

This is for a review. I just want to grasp these questions better. If it helps its one of these. 0.5 mm 5.4 mm 540 mm 5,400 mm

OpenStudy (anonymous):

OF that queston your and my answer are same.